An architectural firm has three types of clients: G = green/new: EC = experienced with a complex project, and ES = ex
Instructions: Round Implied utilization (%) answers to 1 decimal place and Actual and maximum flow rates answers to 3 dec
Answer #1

Solution:

Implied\ Utilisation = \frac{Total\ Demand\ Inflow}{Total\ Capacity}

Hence,

a)

- Implied Utilisation of Needs,

Total Demand = 0.25/day
Total Capacity = 1/7 per day.

Implied\ Utilisation = \frac{0.25}{\frac{1}{7}} = 1.75 = 175\%

Hence, the implied utilization of Needs is 175%

- Implied Utilisation of Design,

Total Demand = 0.40/day (sum of all 3 demands 0.25+0.10+0.05)
Total Capacity = 3/50 per day.

Implied\ Utilisation = \frac{0.40}{\frac{3}{50}} = 6.66 = 666.66\%

Hence, the implied utilization of Design is 666.66%

- Implied Utilisation of Engineering,

Total Demand = 0.35/day
Total Capacity = 0.2/day

Implied\ Utilisation = \frac{0.35}{0.2} = 1.75 = 175\%

Hence, the implied utilization of Engineering is 175%

b)

- Needs:

Actual Flow = 0.25/day
Maximum Flow = 0.25/day

**Can solve only first 4 parts as per Chegg policy**